IPP Progress Report for the week 2011.06.06 - 2011.06.10

(Up to IPP Progress Reports)

Eugene Magnier

I have been focusing much of my time on science related efforts. I tweaked the megacam astrometry parameters and re-ran the u-band MD data. I now have stacks for most MD fields, and photometry of the stacks as well as the individual images (PSF only). I also finished much of the L-dwarf paper analysis: I went through the current common list of L dwarfs and re-examined the photometry to understand and fix if possible the unusual outliers. The figures for this paper are almost all in place now, and the conclusions are also clear. We can now move forward on the paper text. In terms of IPP issues, I spent much of my time on organization. I cleaned up and backed up the dvo catdirs on the manoa cluster, and generally worked to clear out space on those machines. This is needed so we can re-start the rsync jobs, bringing the mini-dvodbs to manoa for distribution. I also put together a plan for upgrades / re-furbishment of the manoa IPP cluster (both ops and dev).

Serge Chastel

vacation

Heather Flewelling

  • vacation last week, so this is the progress report for the week before
  • czar 2 days (monday/tuesday) - encountered a few problems with stdscience (it didn't load up the survey tasks for some reason)
  • addstar
    • restarted dvodb merge
    • restarted addstar on 3pi
    • can't rsync minidvodbs on ipp022 - not enough space, Gene is going to clean it up
    • made dvodb of LAP (all but z band)
    • fixed bugs in addtool -definebyquery
  • made a wiki of the current uses for the manoa cluster: http://svn.pan-starrs.ifa.hawaii.edu/trac/ipp/wiki/Development_Cluster_Uses
  • with Gene's help we figured out isp astrometry needs the luminosity match turned off. I haven't rerun my test set yet.
  • M31 jtrp reprocessing
  • off friday

Roy Henderson

  • PSVO
    • fixed issue where plugin queries not available in jar version of software
    • added constraints to numeric plugin parameters: now popup will only accept 'allowed' values
    • add regular expression constraints to string values
    • added some public documentation on the PSPS wiki (leaving developers' notes on the private 'system' wiki)
    • added more plug-ins: frame counts by survey or filter, get frames by filter
    • removed summary panel, as this can now all be handled with plug-ins
    • fixed some minor issues in PSQB (query builder)
    • adventures with Derby: using client/server mode, rather than embedded. can store to local db, pull out from TOPCAT etc. All hidden from the user.
    • tested new version of TOPCAT
  • Retrieved LAP data for Jim
  • Czar on Thurs/Fri
  • lots of mailing list action

Bill Sweeney

  • Spent the entire week restructuring ipp/Ohana moving the dvo code that performs queries (for example avextact) out of the dvo program and into the dvo library. This will enable this code to be used in other programs.
  • Modified relastro -high-speed use this code to accept variable query strings. This will allow proper motion candidates to be selected using various cuts instead of the hard coded comparisions to 2mass. There is still a bit of cleanup to do on these changes before they can be made available to users.
  • Fixed various processing faults.

Chris Waters

  • Stacking: identified problem with rejection in the convolved stacks as being caused by the large fraction of the total variance contributed by the systematic variance (defined as 0.1 * data). This dominates for inputs with large bright artifacts, and causes faint inputs to be dropped, as they do not have a large systematic variance. Working on bugfixes to re-implement the systematic variance to be defined by the distribution of input values provided by the KMM test. This should provide a better result, and ensure that all inputs are treated equally.
  • asinh scaling: writing phase largely complete, but have not yet tested reading these images back in.
  • Registration: fixed summitExp.dateobs/rawExp.dateobs bug in the registration code on the trunk.
  • Disk space: queued a large amount of unneeded LAP data for cleanup, and continued running ROC repair script.